How many kmols are in 890 cg of sodium phosphate?
1
Expert's answer
2018-01-18T13:49:07-0500
n(Na3PO4)=m(Na3PO4)/M(Na3PO4),
M(Na3PO4)=164 kg/kmols
n(Na3PO4)=890/164=5,427 kmols of sodium phosphate.